# Performance Portable Monte Carlo Neutron Transport in MCDC via Numba

### v1

@@ 2 Performance
To compare MC/DC to another Monte Carlo neutronics software package, we use a time-dependent version of the Kobayashi dog-leg void-duct problem [10, 5]. Figure 4 at left shows the void duct and the location of the neutron source at the opening of the duct. The initial condition is zero flux everywhere. Radiation moves down the void quickly and then penetrates into the walls of the problem. Figure 4 at right shows the duct clearly with the scalar flux solution in a time-averaging window of 10 s at somepoint time from MC/DC and OpenMC [7] (an open-source code written in C++) .
Refer to caption           Refer to caption
@@
Figure 4: Left: Kobayashi problem schematic. Right: Time and space averaged scalar flux solution to the Kobayashi problem run with $1\times 10^{9}$ particle histories from MC/DC (top) and OpenMC (bottom).
Figure 5 at left shows the runtime results of MC/DC compared with OpenMC on a single node of the Quartz supercomputer (two socket 18-core Intel Xeon E5-2695 v4) at Lawrence Livermore National Laboratory (LLNL). Caching in MC/DC was enabled for these results, meaning there is no JIT compilation overhead. We examine the impact of a variance reduction technique called implicit capture [8], which is supported in both codes. OpenMC is about 20 times faster than MC/DC for this problem with implicit capture. Without implicit capture that performance gap falls to 12 times. This performance gap is due to OpenMC’s tracking algorithm, which does not stop particles at mesh crossing events as MC/DC currently does. OpenMC is currently limited to the use of the collision estimator to estimate the scalar flux spatial distribution in time-dependent simulation, whereas MC/DC uses the track length estimator [8], which is computationally more expensive but yields better statistics. This is shown in figure 4 at bottom right where OpenMC’s result has tally cells with a poorly converged solution (displaying as yellow spots) and MC/DC’s (top right) doesn’t. Work is ongoing to implement this particle stop command. When this is enabled, we suspect the performance gaps for this problem will close substantially between MC/DC and OpenMC.
Refer to caption Refer to caption
Figure 5: Left: Wall clock runtime of the Kobyashi problem over particle counts with and without implicit capture (IC) between cached MC/DC (black) and OpenMC (red). Right: Wall clock runtime of the Kobyashi problem over particle counts in uncached MC/DC CPU (black) and GPU (green) modes.
MC/DC is the only open-source Monte Carlo neutron transport code that can perform fully transient analysis for systems like the Kobyashi problem on GPUs. Figure 5 at right shows the wall clock runtime results of the same Kobyashi problem on a single Tesla V100 GPU on the Lassen supercomputer at LLNL as compared to MC/DC on 36 CPU cores (this time with caching disabled) on LLNL’s Quartz supercomputer across particle counts and with/without implicit capture. The JIT compilation overhead incurs a constant runtime before enough work saturates the GPU architecture. For the Kobyashi problem with implicit capture, the GPU versions experience a 21 times speed up with saturation occurring around $10^{6}$ particles. For this problem without implicit capture, the GPU remains unsaturated up to $10^{9}$ but incurs a maximum observed speedup of 24 times over on CPU. This speedup of a whole CPU node to a single GPU compares well to the performance of other production Monte Carlo neutron transport applications [11].
3 Discussions, Conclusions, and Future Work
@@
Monte Carlo/Dynamic Code is a single-source, single-language, single-compiler Monte Carlo neutron transport code that targets modern HPC architectures with CPUs and GPUs. Our performance results demonstrate that MC/DC’s structure (using a Python + Numba JIT + MPI scheme) can produce similar performance to other Monte Carlo neutron transport solvers. While MC/DC currently performs somewhat worse than other production code, we believe this is primarily due to algorithmic differences rather than architectural ones. MC/DC currently supports many novel transient transport methods and variance reduction techniques that are not supported in any other Monte Carlo neutron transport application.
Numba has been found to have performance deficiencies as compared to other high-level languages and portability frameworks in some workloads [12]. We initially explored other high-level language portability frameworks including PyCUDA and PyKokkos and found that for a Monte Carlo neutron test-bed Numba’s performance was in-line with the others considered [13]. Numba has been described as being as difficult to develop in as a traditional low-level language for complex algorithms [14]. We agree with those findings that it can be as difficult—or even harder when debugging—as developing compiled code, but since we are primarily using Numba as a Python-based portability framework, any ease-of-development advantages are only added benefits. MC/DC’s software engineering approach has the added benefit of unifying our glue language and kernel production language.
Work in MC/DC is ongoing. We are continually exploring novel variance reduction techniques and adding new functionality. We will implement a no-stop mesh-boundary tracking algorithm, optimize our data structures, and explore methods of kernel profiling with the goal of identifying code hotspots sections of MC/DC for targeted optimizations. We also plan on adding better tracking schemes to OpenMC to improve it’s solution for transient problems as well. On GPUs specifically, work is ongoing to support compilation to AMD GPUs via a patch to Numba that allows for AMD compilation and execution [^4] https://github.com/ROCm/numba-hip. Intel also offers a patch to compile to their GPUs via Numba [^5] https://github.com/IntelPython/numba-dpex and future work include building support for their accelerators as well. On all GPU types we also plan to extend support to multi-GPU simulations via existing MPI frameworks. We will continue to improve MC/DC, making it a portable application for rapid methods development enabled by Python and Numba.

### v2

@@ 4 Performance
Figure 4: Kobayashi problem schematic.
@@
Figure 5: Time and space averaged scalar flux solution to the Kobayashi problem run with $1\times 10^{9}$ particle histories at various points in time.
To examine the performance of MC/DC we use a time-dependent version of the one-group Kobayashi dog-leg void-duct problem [11, 5]. Figure 4 shows the void duct and the location of the neutron source at the opening of the duct. The initial condition is zero flux everywhere. Radiation quickly moves through the void and then penetrates the walls of the problem, slowly dissipating through time. Figure 5 shows the duct clearly with the scalar flux solution at various points in time.
We solved the Kobyashi problem on HPC systems available at Lawrence Livermore National Laboratory (LLNL): the Dane and Lassen machines. Dane is a CPU-only system with dual-socket Intel Xeon Sapphire Rapids CPUs, each with 56 cores for a total of 112 per node. Lassen has four Nvidia Tesla V100s and two IBM Power 9 CPUs per node. To contrast MC/DC on the CPU against a traditionally developed and compiled code, we will compare performance to another Monte Carlo neutron transport code, OpenMC [^11] https://github.com/openmc-dev/openmc [8] (an open-source code written in C++). We added time-dependent functionality to OpenMC [^12] https://github.com/CEMeNT-PSAAP/openmc/tree/transient so that the same algorithm is implemented in both codes for the Kobyashi problem.
[…]
Figure 6: Left: Wall-clock runtime of the Kobyashi problem over particle counts. Right: Weak scaling efficiency as a function of node count for the Kobyashi problem on Dane (CPU) and Lassen (GPU).
5 Discussion, Conclusions, and Future Work
@@
Monte Carlo/Dynamic Code (MC/DC) is a Monte Carlo neutron transport code that targets modern HPC architectures with CPUs and GPUs. Our performance results demonstrate that MC/DC’s structure using a Python + Numba + MPI + Harmonize scheme can produce similar performance to other Monte Carlo neutron transport solvers. After JIT compilation overhead, MC/DC performs similarly to traditionally compiled production code on a single node for a transient problem of interest. MC/DC exhibits similar weak scaling on CPUs and superior weak scaling on GPUs up to 256 nodes of a given HPC, compared with a CPU-only production code.
Developing using Numba for CPU targets can be as difficult as developing in low-level languages for the complicated algorithms we implement. This agrees with previously published analysis [12]. We found developing the necessary time-dependent features to model the Kobyashi problem in OpenMC to be about as difficult as making changes within MC/DC. For our application, anything gained when using a high-level-language is lost in time and effort spent circumventing unsupported operations and debugging. However, the implementation in OpenMC remains CPU-only, while for MC/DC it took little effort to go from a working CPU implementation to something operating and highly-performing on GPUs. Of course, we use our own specialized event-scheduling library to do this—but Numba allows us to construct a Python-based portability framework fit to our numerical method with the added benefit of unifying our high-level glue language and kernel-production language.
Over the duration of developing MC/DC (starting in 2021) we have seen many improvements to Numba. Compiler error reporting continues to improve (especially in Numba versions 0.59.0$+$), the number of supported operations have grown, and Numba has been extended to additional accelerators like AMD and Intel [^13] https://github.com/IntelPython/numba-dpex GPUs. We have found that the Numba development team fosters a supportive community that is approachable and responsive to questions, comments, and concerns. We believe that as Numba matures we will continue to see performance and development improvements.
Work in MC/DC is ongoing. We are continually exploring novel variance reduction and hybrid Monte Carlo techniques, and adding new functionality. For GPU development specifically we are currently investigating use of unified memory between the CPU and GPU as well as extending support to Intel GPUs. We will continue to improve MC/DC, making it a portable application for rapid methods development enabled by Python and Numba.
@@ 6 Acknowledgments
The authors thank the Numba development team for support using the Numba compiler as well as Damon McDougall and Dominic Etienne Charrier from Advanced Micro Devices for support using Numba-HIP and ROCm compilers. The authors thank the high performance computing staff at Lawrence Livermore National Laboratory for continued support using the Dane and Lassen machines.